OpenStudy (anonymous):

Graph using the x and y intercept method. (remember to plug in 0) Write each intercept as a separate ordered pair using 0. 6x – 3y = -18......where do i start with this?

OpenStudy (anonymous):

i figured out y=2x+6 part but not sure what to do now

OpenStudy (anonymous):

You do not need to solve for y or put into slope-intercept form by any other method. Use constant-over-coefficient.

OpenStudy (anonymous):

i am not sure what that means.

OpenStudy (anonymous):

-18 is the constant 6 is the coefficient of x -3 is the coefficient of y

OpenStudy (gitface):

You are asked to graph. Begin by finding your y-intercept. That is; the value of y when x=0. It also seems the method you are asked to use is to also find hte x intercept, that is when y=0. Begin there.

OpenStudy (anonymous):

'constant-over-coefficient' -18 is the constant 6 is the coefficient of x -18/6 = x-intercept. -3 is the coefficient of y -18/-3 = y-intercept.

OpenStudy (anonymous):

okay thank you gitface! the way you explain stuff makes sense to me, you should be a teacher haha

OpenStudy (anonymous):

sooo then the ordered pairs would be (0,-6) and (3,0) and thats what I graph correct?

OpenStudy (gitface):

The negatives are in the wrong coordinates. (0,6) and (-3,0) are your coordinates.

OpenStudy (anonymous):

oh i see my mistake, thanks!

OpenStudy (gitface):

You're welcome.

OpenStudy (anonymous):

So, you couldn't see that -18/6 = -3 and =18/-3 =6?

OpenStudy (gitface):

That's actually pretty neat. I had never been taught that.

OpenStudy (anonymous):

not saying that, but thats not how i was shown so that was a confusing way of putting it for me. i suck at slopes and i need it dumbed way down!!!! haha

OpenStudy (anonymous):

When you sub in y=0 to find x-intercept, and sub in y=0 to find x-intercept, it is the same as doing constant-over-coefficient, but at the same time.

OpenStudy (anonymous):

i am going to use that in the future because its much faster, thank you! my final exam is tomorrow and that will help a lot! appreciate it!!!